OpenStudy (anonymous):

Fimd a quadratic equation with the following roots-1+4i and -1-4i

OpenStudy (anonymous):

{x-(-1+4i)}{x-(-1-4i)}=0 This implies {(x+1)-4i}{(x+1)+4i}=0 This implies (x+1)^2-(4i)^2=0 This implies (x^2+2x+1)+16=0 This implies x^2+2x+17=0

OpenStudy (anonymous):

@Amberlisa22